# Data-Driven Distributed Optimization via Aggregative Tracking and Deep-Learning

### v2

@@ D-D Proof of Lemma 3
The proof relies on Theorem 2 (cf. Appendix A-A). Hence, we need to check that Assumptions 789, and 10 are satisfied. First, Assumption 7 is guaranteed by the periodicity of $\mathrm{d}^{k}_{x,i}$ and $\mathrm{d}^{k}_{\sigma,i}$ (cf. Assumption 5). Second, Assumption 8 (i.e., the fact that the origin is an equilibrium point of both (25) and the averaged system (50)) is satisfied in light of (6c) in Assumption 6 and since $\theta^{\star}$ is the solution to (4) with $k=\mathcal{K}$ (cf. Assumption 5). Third, the Lipschitz properties in Assumption 9 follow by Assumptions 5 and 6. Fourth, the periodicity of $\mathrm{d}^{k}_{x,i}$ and $\mathrm{d}^{k}_{\sigma,i}$ (cf. Assumption 6) allows for satisfying Assumption 10 with $\nu\left(T\right)\coloneq\frac{L}{1+T}$ where $L>0$ depends on the Lipschitz constants of $\nabla_{3}\ell$ and $\nabla_{3}\tilde{\ell}_{\mathcal{K}}$. Finally, Lemma 2 proves that the origin is an exponentially stable equilibrium of the averaged system (50). Hence, by Theorem 2, there exist $\bar{\gamma}_{1}\in(0,\bar{\gamma}_{2})$ and $r>0$ such that, for all $\gamma\in(0,\bar{\gamma}_{1})$ and $\chi^{0}\in\mathcal{X}$, the origin is an exponentially stable equilibrium of (25). The proof follows by the Converse Lyapunov Theorem (see, e.g., [38, Th. 2.2.1]).
@@ E-E Proof of Lemma 4
The proof is based on timescale separation. First, Lemma 1 provides a Lyapunov function proving the global exponential stability of the origin of (23), i.e., the boundary-layer system associated to (18). Second, Lemma 3 provides a Lyapunov function proving exponential stability of the origin of (28), i.e., the reduced system associated to (18). Third, the dynamics of system (18) and the equilibrium map (19) are Lipschitz continuous (see Assumptions 245, and 6). We thus guarantee exponential stability of $(\theta^{\star},x^{\star},h(\theta^{\star},x^{\star}))$ for system (18) by slightly extending [40, Th. II.3]. In detail, the mentioned result provides global results but requires global results in both the boundary-layer and reduced system, while Lemma 3 (i.e., the stability result of the reduced system) holds only locally. Hence, we conclude that the origin is an exponentially stable equilibrium point of system (18) for any initial condition $(\tilde{\theta}^{0},\tilde{\mathrm{x}}^{0},\mathrm{w}_{\perp}^{0},\mathrm{z}_{\perp}^{0})\in D$. The proof follows by the Converse Lyapunov Theorem, see, e.g., [38, Th. 2.2.1].

### v3

@@ D-D Proof of Lemma 3
The proof relies on Theorem 2 (cf. Appendix A-A). Hence, we need to check that Assumptions 789, and 10 are satisfied. First, Assumption 7 is guaranteed by the periodicity of $\mathrm{d}^{k}_{x,i}$ and $\mathrm{d}^{k}_{\sigma,i}$ (cf. Assumption 5). Second, Assumption 8 (i.e., the fact that the origin is an equilibrium point of both (25) and the averaged system (53)) is satisfied in light of (6c) in Assumption 6 and since $\theta^{\star}$ is the solution to (4) with $k=\mathcal{K}$ (cf. Assumption 5). Third, the Lipschitz properties in Assumption 9 follow by Assumptions 5 and 6. Fourth, the periodicity of $\mathrm{d}^{k}_{x,i}$ and $\mathrm{d}^{k}_{\sigma,i}$ (cf. Assumption 6) allows for satisfying Assumption 10 with $\nu\left(T\right)\coloneq\frac{L}{1+T}$ where $L>0$ depends on the Lipschitz constants of $\nabla_{3}\ell$ and $\nabla_{3}\tilde{\ell}_{\mathcal{K}}$. Finally, Lemma 2 proves that the origin is an exponentially stable equilibrium of the averaged system (53). Hence, by Theorem 2, there exist $\bar{\gamma}_{1}\in(0,\bar{\gamma}_{2})$ and $r>0$ such that, for all $\gamma\in(0,\bar{\gamma}_{1})$ and $\chi^{0}\in\mathcal{X}$, the origin is an exponentially stable equilibrium of (25). The proof follows by the Converse Lyapunov Theorem (see, e.g., [1, Th. 2.2.1]).
@@ E-E Proof of Lemma 4
The proof is based on timescale separation. First, Lemma 1 provides a Lyapunov function proving the global exponential stability of the origin of (23), i.e., the boundary-layer system associated to (18). Second, Lemma 3 provides a Lyapunov function proving exponential stability of the origin of (28), i.e., the reduced system associated to (18). Third, the dynamics of system (18) and the equilibrium map (19) are Lipschitz continuous (see Assumptions 245, and 6). We thus guarantee exponential stability of $(\theta^{\star},x^{\star},h(\theta^{\star},x^{\star}))$ for system (18) by slightly extending [4, Th. II.3]. In detail, the mentioned result provides global results but requires global results in both the boundary-layer and reduced system, while Lemma 3 (i.e., the stability result of the reduced system) holds only locally. Hence, we conclude that the origin is an exponentially stable equilibrium point of system (18) for any initial condition $(\tilde{\theta}^{0},\tilde{\mathrm{x}}^{0},\mathrm{w}_{\perp}^{0},\mathrm{z}_{\perp}^{0})\in D$. The proof follows by the Converse Lyapunov Theorem, see, e.g., [1, Th. 2.2.1].